OpenStudy (anonymous):

Use integration by parts to prove the following reduction reduction formulae.

OpenStudy (anonymous):

\[\int\limits_{}^{}x^{n}\ln(x) dx = \frac{ x^{n+1} }{ (n+a)^{2} }[-1+(n+1)lnx]+c\]

OpenStudy (anonymous):

i know ill be using \[\int\limits_{}^{}udv=uv-\int\limits_{}^{}v du\]

OpenStudy (anonymous):

and so far i've let u=x^(n+1)/(n+1)^2 so du=x^n/n+1 and i've let dv=dx so v=x

OpenStudy (ash2326):

Is that "a" on the right side?

OpenStudy (anonymous):

no it's x^n

OpenStudy (anonymous):

oh it should be n+1 not n+a

OpenStudy (ash2326):

You should choose dv as x^n

OpenStudy (anonymous):

no a's in the equation at all sorry.

OpenStudy (anonymous):

and should u be what?

OpenStudy (ash2326):

\[\int x^n \ln x dx=\frac{x^{n+1}}{n+1}\times \ln x -\int \frac d {dx} \ln x\times \frac{x^n}{n+1} dx\] \[\frac{x^{n+1}}{n+1}\times \ln x -\int \frac 1 {x} \times \frac{x^n}{n+1} dx\] \[\frac{x^{n+1}}{n+1}\times \ln x -\int \frac{x^{n-1}}{n+1} dx\] Can you solve from hee?

OpenStudy (anonymous):

so i let u=ln x so du=1/x and i let dv = x^n so v=nx^(n-1) ??

OpenStudy (ash2326):

*here?

OpenStudy (anonymous):

was my previous statement right?

OpenStudy (anonymous):

and im having problem solving the integral part.

OpenStudy (ksaimouli):

use tabular integration

OpenStudy (anonymous):

saimoulis meathod is easy tabular integration is much eaiser